OpenStudy (anonymous):

Find the area of the region under the curve f(x)=tanx on [0,π/2].

OpenStudy (anonymous):

its infinity

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2}\tan x dx=\ln|secx||_{0}^{\pi/2}=\ln|\infty |-0=\infty\]

OpenStudy (anonymous):

im unclear on how you got this answer, can you tell me the steps you took?

OpenStudy (anonymous):

i just integrated tanx in the limit 0 to pi/2

OpenStudy (anonymous):

so tyhe integration of tan x is ln|secx| ??

OpenStudy (anonymous):

yes..don't u know that??

OpenStudy (anonymous):

i thought the integral of tan was sec^2

OpenStudy (anonymous):

no the derivative is sec^2

OpenStudy (anonymous):

oooh, oh ok. and then how did you get this..ln|∞|−0=∞ ???

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

how did you get... ln|∞|−0=∞ ???